Calculate the equilibrium concentrations of C2HsCOO and C2HsCOOH in a solution prepare by dissolving 0.04 mole of propionic acid in 150 ml of water.
Answer #1

molarity of C2H5COOH = no of moles/volume in L

                                          =0.04/0.15 = 0.267M

                      C2H5COOH (aq)--------------> C2H5COO^- (aq) + H^+(aq)

              I        0.267                                          0                             0

              C         -x                                             +x                            +x

              E       0.267-x                                        +x                           +x

                            Ka          = [C2H5COO^-][H^+]/[C2H5COOH]

                              1.34*10^-5   = x*x/0.267-x

                           1.34*10^-5 *(0.267-x) = x^2

                                          x = 0.0019

                      [C2H5COO^-]   = x = 0.0019M

                          [H^+]              = x   = 0.0019M

                     [C2H5COOH]    = 0.267-x   = 0.267-0.0019    = 0.265M